# Hybrid Quantum Solvers in Production: how to succeed in the NISQ era?

### v6

@@ Abstract
Index Terms:
Hybrid Quantum-Classical Computing, Combinatorial Optimization, Quantagonia, D-Wave, Quantum Computing.
I Introduction
Recent years have seen significant progress in quantum computing (QC), primarily as a result of the fast development of the technology and the advances made in the democratization of its access [1]. As a result, QC has already helped several application domains launch various proofs of concept. Frequently, these endeavors depend on hybrid systems, which symbolize the immediate future of this field. This is so since hardware capabilities play a major role in the adoption of quantum approaches to solve industrial use cases. Being immersed in the noisy intermediate scale quantum (NISQ, [2]), today’s quantum technology is neither entirely dependable nor competent to resolve medium- or large-sized problems. In this context, the challenge lies in finding the way to build fruitful harmony between classical and quantum computing, assembling a tandem that, eventually, outperforms previous purely classical conceptualizations.
@@ I Introduction
      According to the layout established by the workflow of the hybrid solver, following the guidelines of [4]. This means focusing on the logical arrangement of the quantum and classical modules within the method. Note that this work focuses on giving an overview of the hybrid solver strategies with emphasis on the designable-by-designer steps. This means that the taxonomy will not include any consideration with respect to vertical hybrid workflows, i.e. those with “controlling activities required to control and operate” either a classical or a quantum subroutine, as these operations are not key functional stages of the algorithmic scheme of the solver.
@@
  • Given the previous theoretical framework, gaining insight into the performance of those hybridization strategies nowadays in production and close to industrial markets. To do that, an experimentation on two hybrid solvers is conducted: the LeapHybridBQMSampler (LeapBQM) contained in D-Wave’s Hybrid Solver Service (HSS, [6]), and Quantagonia’s Hybrid Solver [^1] https://www.quantagonia.com/hybridsolver. For these tests, the QOPTLib benchmark [7] has been employed, which is composed of 40 instances equally distributed over four combinatorial optimization problems: the Traveling Salesman Problem (TSP, [8]), the Vehicle Routing Problem (VRP, [9]), the one-dimensional Bin Packing Problem (BPP, [10]), and the Maximum Cut Problem (MCP, [11]).
Figure 1: Classification of HS schemes. Image taken from [5].
The rest of this paper is structured as follows: In the following Section II, we perform the classification of five of the most renowned and used hybrid solvers to summarize current trends for hybrid solver designs, which answers the main question posed in the title of the section: “Where might the quantum advantage be?”. After that, Section III describes the experimentation conducted with LeapBQM and QHS, including the experimental setup and results. Finally, Section IV closes this document by drawing some conclusions and outlining future research lines.
II Classifying hybrid solvers: Where might the quantum advantage be?
@@
The following are examples of well-known micro hybrid split schemes:
@@
  • QBSolv: QBSolv is a hybrid method developed by D-Wave, which can be described as “a decomposing solver for solving a QUBO problem by splitting it into pieces [^2] https://github.com/dwavesystems/qbsolv. The decomposing procedure of this method starts with an initial solution provided by a classical Tabu Search algorithm [20]. A quantum method later improves that solution by sending a user-specified fraction of it to the QPU. This subproblem includes the most energetic variables, of size fraction, of the initial problem. In practice, this value is usually between 0.05 and 0.15.
@@
  • Quantagonia’s Hybrid Solver: established in November 2021, Quantagonia’s main objective is to bridging the gap between quantum computing, today mainly in lab, and real industry needs. Among the portfolio of products offered by this German company, QHS is a hardware-agnostic method for addressing Linear Programming, Mixed-Integer Programming and QUBO problems. Despite the fact that QHS has not been employed in any research paper yet, it is currently being used by leading companies such as Strangeworks, MathPlan, T-Systems, and Adesso, which reinforces its practical application.
@@
Figure 6: Quantagonia’s Hybrid Solver workflow. CH = classical heuristic.
QM = Quantum Module. B&B = Branch and Bound algorithm.
Lastly, we depict in Table I a summary of the complete categorization described along this section.
TABLE I: Categorization of the hybrid solvers described
@@
QBSolv Cooperative - Imbrication Exploration: classical computing, Exploitation: QPU,
@@
HSS Cooperative - Imbrication (each generated thread) Exploration: classical computing, Exploitation: QPU,
QHS Cooperative - Imbrication Exploration: QPU, Exploitation: classical computing,
III Experimentation
@@
    LeapBQM, and QHS have successfully demonstrated to be near the real industry, being currently used by some leading companies of different types. Delving into advanced solvers such as these provides the reader with an approximate idea of the performance and philosophy of methods that are currently deployed in real production.
@@
Thus, Table II shows the results obtained after 10 independent runs per instance. For each instance, we represent the result average obtained for each solver, as well as the standard deviation and the median value. Note that all these values correspond to the original objective functions above described. Aiming to enhance the replicability of this work, all the 40 instances used and the results obtained are openly available in [33].
TABLE II: Results obtained by LeapBQM and QHS for the whole QOptLib benchmark. Note that the values correspond to the original objective function already especified in Section III. Best average values per instance have been highlighted in bold.
@@
As a general conclusion, we can see through the results depicted in Table II that QHS clearly outperforms LeapBQM. While for the MCP both methods perform equally well, finding the optimum in all runs and for the entire benchmark, this trend is not replicable for the rest of the problems. For TSP, BPP, and VRP, LeapBQM struggles to scale as the size of the problem increases. This is not the case for QHS, whose scalability limits are higher than the ones studied in this paper for the TSP and BPP and whose performance is acceptable for the whole VRP benchmark. In fact, QHS has found the optimum value in all the executions for all the TSP, BPP, and MCP instances, as well as for these VRP cases comprised of less than seven nodes.
IV Conclusions & Further Work
@@
This change in algorithmic design may be one of the keys to the preeminence of QHS versus LeapBQM, and should make the reader reflect on the wide variety of possibilities when approaching the development of a hybrid algorithm. Thus, we encourage researchers involved in this field to devise novel hybridization mechanisms that maximize the synergies between the two computational paradigms. Undoubtedly, the optimal placement of the puzzle pieces will unlock the secret to success in the NISQ era.

### v7

@@ Abstract
Keywords:
Hybrid Quantum-Classical Computing, Combinatorial Optimization, Quantagonia, D-Wave, Quantum Computing
1 Introduction
Recent years have seen significant progress in quantum computing (QC), primarily as a result of the fast development of the technology and the advances made in the democratization of its access [24]. As a result, QC has already helped several application domains launch various proofs of concept. Frequently, these endeavors depend on hybrid systems, which symbolize the immediate future of this field. This is so since hardware capabilities play a major role in the adoption of quantum approaches to solve industrial use cases. Being immersed in the noisy intermediate scale quantum (NISQ, [23]) era, today’s quantum technology is neither entirely dependable nor competent to resolve medium- or large-sized problems. In this context, the challenge lies in finding the way to build fruitful harmony between classical and quantum computing, assembling a tandem that, eventually, outperforms previous purely classical conceptualizations.
@@ 1 Introduction
      According to the layout established by the workflow of the hybrid solver, vertical versus horizontal, following the guidelines of [21]. This means focusing on the logical arrangement of the quantum and classical modules within the method. Note that this work focuses on giving an overview of the hybrid solver strategies with emphasis on the designable-by-designer steps. This means that the taxonomy will not include any consideration with respect to vertical hybrid workflows, i.e. those with “controlling activities required to control and operate” either a classical or a quantum subroutine, as these operations are not key functional stages of the algorithmic scheme of the solver.
@@
  • Given the previous theoretical framework, gaining insight into the performance of those hybridization strategies nowadays in production and close to industrial markets. To do that, an experimentation on two hybrid solvers is conducted: the LeapHybridBQMSampler (LeapBQM) contained in D-Wave’s Hybrid Solver Service (HSS, [7]), and Quantagonia’s Hybrid Solver [^1] https://www.quantagonia.com/hybridsolver. For these tests, the QOPTLib benchmark [17] has been employed, which is composed of 40 instances equally distributed over four combinatorial optimization problems: the Traveling Salesman Problem (TSP), the Vehicle Routing Problem (VRP), the one-dimensional Bin Packing Problem (BPP), and the Maximum Cut Problem (MCP).
Figure 1: Classification of HS schemes. Image based on [30].
The rest of this paper is structured as follows: In the following Section 2, we perform the classification of five of the most renowned and used hybrid solvers to summarize current trends for hybrid solver designs, which answers the main question posed in the title of the section: “Where might the quantum advantage be?”. After that, Section 3 describes the experimentation conducted with LeapBQM and QHS, including the experimental setup and results. Finally, Section 4 closes this document by drawing some conclusions and outlining future research lines.
2 Classifying hybrid solvers: Where might the quantum advantage be?
@@
The following are examples of well-known micro hybrid split schemes, that is, solvers that deal with a single activity, in which some operations are quantum and others are classical (in an interative fashion):
@@
  • QBSolv: QBSolv is a hybrid method developed by D-Wave, which can be described as “a decomposing solver for solving a QUBO problem by splitting it into pieces [^2] https://github.com/dwavesystems/qbsolv. The decomposing procedure of this method starts with an initial solution provided by a classical Tabu Search algorithm [13]. A quantum method later improves that solution by sending a user-specified fraction of it to the QPU. This subproblem includes the fraction of most energetic variables of the initial problem. In practice, this value is usually between 0.05 and 0.15.
@@
  • Quantagonia’s Hybrid Solver: established in November 2021, Quantagonia’s main objective is to bridge the gap between quantum computing, mainly in labs, and real-world industry needs. Among the portfolio of products offered by this German company, QHS is a hardware-agnostic method for addressing Linear Programming, Mixed-Integer Programming and QUBO problems. Despite the fact that QHS has not been employed in any research paper yet, it is currently being used by leading companies such as Strangeworks, MathPlan, T-Systems, and Adesso, which reinforces its practical application.
@@
Figure 6: Quantagonia’s Hybrid Solver workflow. CH = classical heuristic. QM = Quantum Module. B&B = Branch and Bound algorithm.
Lastly, we depict in Table 1 a summary of the complete categorization described along this section.
Table 1: Categorization of the hybrid solvers described
@@
QBSolv Cooperative - Imbrication Exploration: classical computing, Exploitation: QPU
@@
HSS Cooperative - Imbrication (each generated thread) Exploration: classical computing, Exploitation: QPU
QHS Cooperative - Imbrication Exploration: QPU, Exploitation: classical computing
3 Experimentation
@@
    LeapBQM and QHS have successfully demonstrated to be near the applied industry, being currently used by some leading companies of different types. Delving into advanced solvers such as these provides the reader with an approximate idea of the performance and philosophy of methods that are currently deployed in real production.
@@
Thus, Table 2 shows the results obtained after 10 independent runs per instance. For each instance, we represent the result average obtained for each solver, as well as the standard deviation and the median value. Note that all these values correspond to the original objective functions above described. Aiming to enhance the replicability of this work, all the 40 instances used and the results obtained are openly available in [18].
Table 2: Average, Standard Deviation and Median of the results obtained by LeapBQM and QHS for the whole QOptLib benchmark. Note that the values correspond to the original objective function already especified in Section 3. Best average values per instance have been highlighted in bold.
@@
As a general conclusion, we can see through the results depicted in Table 2 that QHS clearly outperforms LeapBQM. While for the MCP both methods perform equally well, finding the optimum in all runs and for the entire benchmark, this trend is not replicable for the rest of the problems. For TSP, BPP, and VRP, LeapBQM struggles to scale as the size of the problem increases. This is not the case for QHS, whose scalability limits are higher than the ones studied in this paper for the TSP and BPP and whose performance is acceptable for the whole VRP benchmark. In fact, QHS has found the optimum value in all the executions for all the TSP, BPP, and MCP instances, as well as for these VRP cases comprised of less than seven nodes.
4 Conclusions & Further Work
@@
This change in algorithmic design may be one of the keys to the preeminence of QHS versus LeapBQM, and should make the reader reflect on the wide variety of possibilities when approaching the development of a hybrid algorithm. Indeed, algorithmic design is a crucial task not only for methods based on quantum computing [14], but also for classical techniques [22]. Thus, we encourage researchers involved in this field to devise novel hybridization mechanisms that maximize the synergies between the two computational paradigms. Undoubtedly, the optimal placement of the puzzle pieces will unlock the secret to success in the NISQ era, and beyond.
Acknowledgments
This work was supported by the Basque Government through HAZITEK program (Q4_Real, ZE-2022/00033) and through Plan complementario comunicación cuántica (EXP. 2022/01341) (A/20220551). It was also supported by the Spanish CDTI through Misiones Ciencia e Innovación 2021 program (CUCO: Quantum Computing and its Application to Strategic Industries, Grant MIG-20211005).